OpenStudy (anonymous):

For questions 1 through 6 graph each exponential function

OpenStudy (anonymous):

1.f(x) = 6x 2.f(x) =(1/7)^x 3.f(x) = 4^x+2 4. f(x)=(1/3)^x-3 5.Using complete sentences, explain where exponential functions might be applied in the real world. 6.Using complete sentences, explain the limitations on the base of an exponential function.

OpenStudy (anonymous):

I googled number 6 and this came up. I know it probably doesnt matter to you since this was six months ago but to people that google it as I did for help I am putting this here. For questions 1-4 all you have to do is go to geogebra.org and type in the function. It will then graph it for you. You then go to file and select export and can copy it to your document. The teacher is just making sure you know how to put your parenthesis and they you can tell the difference in what the graph looks like.